# https://www.projectaria.com/datasets/nymeria

### v1

@@ Abstract
We introduce Nymeria - a large-scale, diverse, richly annotated human motion dataset collected in the wild with multiple multimodal egocentric devices. The dataset comes with a) full-body 3D motion ground truth; b) egocentric multimodal recordings from Project Aria devices with RGB, grayscale and eye-tracking cameras, IMUs, magnetometer, barometer, and microphones; and c) an additional "observer" device providing a third-person viewpoint. We compute world-aligned 6DoF transformations for all sensors, across devices and capture sessions. The dataset also provides 3D scene point clouds and calibrated eye gaze. We derive a protocol to annotate hierarchical language descriptions of in-context human motion, from fine-grain pose narrations, to atomic actions and coarse activity summarization. To the best of our knowledge, the Nymeria dataset is the world’s largest in-the-wild collection of human motion with natural and diverse activities; first of its kind to provide synchronized and localized multi-device multimodal egocentric data; and the world’s largest dataset of motion with language descriptions. It contains 1200 recordings totalling 300 hours of daily activities of 264 participants across 50 locations, travelling a total of 399Km . The motion-language descriptions provide 310.5K sentences in 8.64M words from a vocabulary size of 6545. To demonstrate the potential of the dataset we define key research tasks for egocentric body tracking, motion synthesis, and action recognition and evaluate several state-of-the-art baseline algorithms. Data and code will be open-sourced.
@@
motion capture egocentric multimodal large dataset
@@ 1 Introduction
The advent of AI is leading to a surge of smart glasses [27, 10, 8, 4, 5, 1, 3, 2] and other AI enabled wearable device applications. These devices not only provide seamless access to LLM-based AI assistants, but also are multi-modal data-capture vehicles that can capture immediate and long-term personalized context, allowing such AI assistants to evolve into the next generation of human-centric, contextualized AI and – long-term – combined with AR/VR technology to unlock a new generation of contextualized computing.
Important context includes the wearer’s own motion (ego-motion) and their actions. Reliably estimating this context solely from body-worn devices is very challenging as typically the sensor measurements alone are insufficient to fully constrain the full body pose – requiring the use of generative methods and techniques from the field of full body motion synthesis. In practice, most research in this direction is severely limited by the existing training data, having to rely on either simulations that lack realism and completeness [85, 40, 41, 11, 12, 47], or limited data collections lacking scale and diversity [34, 66, 45, 90, 46, 44]. There are three key technical challenges:
  • Obtaining full-body motion ground truth. Vision-based solutions such as the ones relying on optical markers [58, 66, 44] or multi-view videos [96, 43, 50, 30] are adversely affected by line-of-sight visibility, and require a complex multi-camera-setup to cover a limited range of motions inside a constrained volume. Inertial-based solutions [6, 9] are an alternative, but suffer from accumulation of error and thus are inferior in global accuracy [78, 59].
  • Multi-device synchronization. Combining multiple capture devices or ground-truth systems requires accurate time- and 6DoF-alignment, which can be challenging when using off-the-shelf solutions that cannot be modified or can lack support for universal protocols. Instead, most existing datasets work around this problem using visual cues [43, 30, 96] or audio [34] signals. However, such approaches are limited in accuracy and reliability. Especially for long recordings, clock-drift accumulates and these methods become intrusive and interrupting to the activity. Consequently, the average motion length is very short for existing datasets (cf. Tab. 2).
  • Data processing and annotations. These are critical for a dataset to develop its full potential. In addition to the raw body motion, multi-device co-localization and 3D scene representation, we believe natural language descriptions are crucial for future research directions. Existing work provides simple descriptions without scene context [71, 31, 23], which is typically done by playing back motion clips for annotators in isolation. Compared to the text corpus for training LLMs [13, 65, 86], the scale of motion-language dataset is significantly smaller.
Seq Qty Ppl Sce Loc Pose Img IMU Gaze Traj Sent Word Voc
@@
Table 1: Unprecedented scale of Nymeria dataset. We capture 1200 sequences of 300 hours daily activity from 264 people, 50 locations with 20 scenarios in the wild. The data provides 260M body poses, 201.2M images, 11.7B IMU samples, 10.8M gazes. The head moving trajectory exceeds 399Km. The motion-language descriptions provide 310.5K sentences in 8.64M words from a vocabulary size of 6545.
To fill the gap and accelerate the research, we introduce the Nymeria dataset: the world largest collection of in-the-wild human motion with rich, multimodal egocentric captures and annotations. Tab. 1 summarizes the core statistics of our dataset. It contains 300 hours with 260M poses of 264 participants across 50 locations. Focusing on daily activities, we designed 20 indoor and outdoor scenarios to capture the diversity and authenticity of how people interact with the environments and others, e.g. playing games/sports with friends, dinning in the cafeteria, working in the office, hiking in the woods, biking on the road etc. All recordings are 15 minutes long, providing ample time for natural activities with spontaneous, unscripted actions and interactions.
We use an inertial-based XSens Mocap [6] suite to capture full-body motion. We use Project Aria [27] glasses in all recordings to collect RGB, grayscale and eye tracking (ET) videos, multichannel audio and, raw data from inertial measurement units (IMUs), magnetometer and barometer. Similar multimodal data from both wrists is collected with a wristband device (miniAria) constructed by repackaging the electronics and sensors from Project Aria glasses to be wrist-worn. A non-intrusive custom hardware solution is used to achieve sub-millisecond time-synchronization across all devices and over the duration of each sessions. We use Project Aria Machine Perception Service (MPS) [7] to obtain accurate 6DoF device trajectories for headset and wristbands, 3D semi-dense point clouds and calibrated eye gaze – and align everything into a single, metric, world coordinate frame. In addition, XSens motion is also registered into the same reference by correcting the drift via an optimization. Overall, the dataset captured 399Km of travel by the participants for a total of 201.2M egocentric images, 11.7B IMU samples and 10.8M gaze points.
The Nymeria dataset also stands out as the world’s largest motion-language dataset. Annotators generated in-context narrations by observing the synchronized playback of rendered egocentric view, third-person view, and human motion from recorded sessions. We designed a hierarchical annotation schema to allow obtaining descriptions at different levels of detail. In total, we annotated 38.6 hours of fine-grained body poses, 207 hours with simplified atomic actions and 196 hours with activity summarization. The total descriptions contain 310.5K sentences and 8.64M words from 6545 vocabulary size.
2 Related Works
dataset q/h p/M $\mu$/m pp nl/K voc pm hd 3p wd ga od gz sr mp hh
@@ 2 Related Works
Table 2: Human motion datasets by releasing date. Columns 2 to 5 show the recorded activity in hour (q/h), body poses in million frames (p/M), average sequence duration in minute ($\mu$/m) and number of participants (pp). We then compare motion with natural language annotations w.r.t. number of descriptions (nl/K) and vocabulary size (voc). The remaining columns cover the following features: using parametric human model and meshes for ground-truth motion (pm), with egocentric head-mounted device (hd), with third-person perspectives (3p), with egocentric wrist-worn device (wd), using global motion positioning (ga), with outdoor scenarios (od), with eye gaze (gz), with 3D scene representations (sr), with multi-people scenarios (mp), and human-human interactions (hh). Note EgoExo4D reports 1422h in the paper, which sums the recording hour by cameras. The total activity is 180h, with 88.8h annotated with MSCOCO keypoints. The underline numbers are reported for the full dataset.
2.1 Human motion datasets
Large-scale motion dataset.
Large-scale datasets, both real and synthetic, provide invaluable 3D annotations for analyzing movements of individuals and groups. AMASS [58] integrates 15 mocap datasets into a unified SMPL [52] format, facilitating extensive motion analysis. 3DPW [59] employs IMUs and mobile cameras to establish the first multi-person, in-the-wild dataset with 3D keypoints. DivaTrack [90] merges synchronized body motion data and IMU signals, supporting a variety of body types and motion classes. Leveraging the GTA-V engine, GTA-Human [15] synthesizes a rich array of 3D-rendered actions, offering crucial supplements to the real data usually gathered indoors. Focusing on interaction, ParaHome [44] and HPS [34] capture human dynamics from egocentric and third-person views. ParaHome [44] records precise home activities and object movements in a unified space, whereas HPS [34] compiles an exhaustive egocentric dataset with 3D scene scans, videos, IMU data, and reconstructions of humans engaging with their surroundings, enriching the landscape of interaction perception and synthesis research.
Egocentric multimodal motion dataset.
To advance first-person perception research, several egocentric video datasets [20, 22, 21, 29] have been developed, focusing on hand-object interaction and action recognition but often omitting 3D ground-truth data. To address this gap, datasets with comprehensive multimodal sensory data and annotations have been introduced to better capture human activities within 3D scenes. UnrealEgo [96] provides naturalistic single-person stereo images and sequences in dynamic 3D environments. EgoBody [96] is the first extensive dataset offering both egocentric and third-person views, with detailed 3D annotations for both individuals and scenes. EgoHumans [43] presents the first collection of multi-human 3D activities from egocentric perspectives. Ego-Exo4D [30], AEA [57, 56] and ADT [66] further enrich the field by documenting skilled or common daily activities.
Motion-language dataset.
Understanding human movement semantics is enhanced by datasets with semantic labels, which vary in annotation depth. Some motion-capture (mocap) datasets [37, 33, 28, 35, 80, 14] offer sparse annotations with limited action categories, while others provide detailed semantic labels: BABEL [72] annotates AMASS [58] sequences with dense, precisely-aligned action labels via crowdsourcing. The KIT Motion-Language Dataset [71] is the first to offer complete natural language descriptions for locomotions. HumanML3D [31] enriches HumanAct12 [33] and AMASS [58] motion clips with three distinct text descriptions per clip. Motion-X [50] stands out as the first comprehensive collection of 3D whole-body motions, including facial expressions and hand gestures, with fine-grained text annotations at both sequence and frame levels.
Tab. 2 shows an overview summary. The existing human motion datasets are either captured without egocentric devices [58, 31, 50, 90, 44, 18], limited in scale [34, 96, 43, 56, 66, 44, 18], lack parametric human model for motion representation [30]. The Nymeria dataset advances the field by improving all axes.
2.2 Egocentric motion understanding
2.2.1 Body tracking and synthesis from motion sensors.
IMU-based MoCap systems are widely adopted for capturing motion in real-world scenarios. While tracking with fewer IMUs is much preferred in practice, it leads to ill-posed body tracking. To tackle the problems of sparsity and noise, recent methods exploit data-driven approaches to learn a motion model from large datasets. DIP [36] uses a prior based on a bi-directional RNN trained on synthesized IMU measurements from AMASS [58]. Transpose [93] further improves the tracking accuracy and addresses root translation estimation. PIP [92] adds a physics layer to refine the joint orientations and root motion predictions. TIP [41] improves root motion estimation by predicting stationary points. Some prior methods tailor tracking systems for specific off-the-shelf IMU-enabled devices, with some utilizing AR/VR devices for the head and hands [40, 39, 90], while others use mobile devices [60]. DiffusionPoser [89] takes the opposite approach by allowing the number of IMUs to be determined during inference using an inpainting technique on a pre-trained diffusion model. Alternatively, combining sparse IMUs with other egocentric modalities can enhance tracking accuracy. EgoLocate [91] combines IMUs with a monocular camera for motion tracking, device localization and mapping. MocapEveryone [46] utilizes two wrist IMUs and a head-mounted camera to capture 3D human motion in diverse environments.
2.2.2 Motion synthesis from text.
Text-conditioned motion synthesis is an active research area, fueled by the availability of annotated human motion datasets such as BABEL [72] and HumanML3D [31]. Early approaches leverage variational autoencoders (VAEs) to generate diverse motions from a textual description [70, 32]. Subsequent works learn to tokenize human motion with a vector quantized variational autoencoder (VQ-VAE) to synthesize motion from text [38, 98]. MotionClip [82] also explores representation learning by training a latent space that aligns with CLIP [73] using a transformer-based autoencoder. With the advent of diffusion models, researchers have shown that higher quality and diversity in text-to-motion synthesis can be achieved with a straightforward training process [83, 79, 19]. Ongoing research refines these these text-conditioned diffusion models by incorporating spatial constraints [42], generating longer motion sequences [83, 99], learning a latent space to improve generation speed and quality [17], and enhancing motion in response to fined-grained textual descriptions [95, 94]. Differing from existing human motion datasets coupled with text descriptions, the Nymeria dataset contains hierarchical language descriptions of diverse scenarios. We believe that the unprecedented diversity provided by our dataset will serve as the key ingredient for inspiring breakthroughs in the field.
3 Nymeria Dataset
Our goal is to record long-duration diverse daily activities with real-world settings. In addition to motion capture and headset, we also consider wristband devices, where wrist-worn sensors closely resemble current AR/VR headsets and are shown to better constrain egocentric body tracking. The remaining section details the capture setup, recording protocols, scenarios, and core statistics.
3.1 Data capture system
@@
Figure 2: Capture setup. From left to righ: (a) a full-dressed participant who wears the mocap suit, Project Aria glasses, two miniAria wristbands and a synchronization device; (b) the set of wearble recording devices; (c) sensor suite of Project Aria and (d) miniAria wristband and how it is worn on the wrist.
@@
The capture setup is illustrated with Fig. 2. To record data, participant wears a mocap suit, a pair of Project Aria glasses, two wristbands, and a synchronization device. We use XSens MVN Link [6] for motion capture, which is a tight-fit body suit wired with 17 inertial trackers and a magnetometer. Compared to other inertial-based systems, MVN Link supports on-device recording, making it ideal for in-the-wild motion capture. To record multimodal egocentric data, we use Project Aria glasses [27] as a lightweight headset. The sensor suite includes 1 RGB camera, 2 grayscale peripheral cameras, 2 ET cameras, 2 IMUs, 1 barometer, 1 magnetometer, 7 microphones, 1 thermometer, GNSS, WiFi and BT. To resemble future wearable devices and AR/VR setup, we repackaged the electronics and sensors into a new wrist-worn device called miniAria. The miniAria wristband matches Project Aria glasses’s sensing ability, with exclusion of microphones, barometer, magnetometer and ET cameras. Appendix 0.A provides further detail on sensor configuration and recording profiles.
@@
Project Aria can record an externally provided time signal to aid synchronization. We further developed the ability for MVN Link to accept the same signal. To time align all devices, a synchronization device is developed to supply the time signal to Project Aria glasses, miniAria wristbands and mocap suit. This device can optionally receive time from a wireless server located in radio range ($\sim$100m), which enables time alignment across multiple Project Aria devices to sub-millisecond accuracy. The synchronization between mocap and Aria is within 1 motion frame i.e. 4.2 ms. To capture multiple people simultaneously, we replicate the described setup per participant and leverage a common time server.
@@
Figure 3: Diverse scenarios by diverse people. We show different participants performing different indoor and outdoor activities at different locations. The top left, bottom left and right views of each subfigure shows the egocentric RGB camera view, the auxillary third-person view and the motion, respectively.
3.2 Recording protocols
3.2.1 Logistics.
We recruit diverse participants, targeting a uniform demographics among gender, age, height, and weight. During recording, an observer wearing Aria glasses is present, who is trained to participants at distance as the moving camera. This third-person holistic view provides significant extra context to improve motion-language narration. Besides participants and observers, two operators are onsite to manage the collection. All people interact naturally as per scenario requires, which lead to rich dynamics as opposed to staged motion. To scale collection, operators are trained to follow a 3-step procedure to complete each recording. It starts with participants performing <1min motion calibration for XSens, then <1min eye calibration for gaze estimation, and last a 15-20min activity. For one-participant collection, each person spent half a day performing 5 scenarios. For two-participant session, people are hired for a full day to collaborate on 8 scenarios. A bulk of data is captured at houses. Each location is rented for 2-4 days.
@@ 3.2 Recording protocols
We define 20 scenarios of common daily activities (cf. Sec. 0.C.1 for definitions). Typical indoor scenarios include objects searching, house touring, cooking, dining, working, entertainment etc, whereas outdoor activities include hiking, biking, sports, gardening etc. To record natural and authentic motions, we debrief participants with high-level guidelines, e.g. “in the following 15-20min, we will get food in the cafeteria and eat in the patio”. Most scenarios encourage human interactions, This also enables operators to prompt in-context activities to increase dynamics. Figure 1 and Fig. 3 visualize examples of the diverse scenarios.
Cauc.HispanicAAEASASEAMixed525252525555151515152525252522222222101010106666525252527777181818181313131321212121222222222222(a) ethnicity18-2425-3030-3535-4040-4545-50222222223737373724242424313131311010101011111111333333332424242426262626222222221313131315151515(b) age group$<$150150-160160-170170-180180-190$>$19044444949494955555555222222225555000044442626262648484848484848487777(c) height/cm40-5050-6060-7070-8080-9090-100272727274848484836363636121212126666666622221111111137373737393939393030303015151515(d) weight/Kg
Figure 4: Demographics by gender. Participants approximately split even among female and male. The ethnicity includes Caucacian (Cauc.), Hipanic, African American (AA), East Asian (EA), South Asian (SA), Southern Eastern Asian (SEA) and mixed.
3.3 Statistics
We collected 1200 recordings from 264 participants, which amounts to 300 hours of activities with 15min per recording on average. The accumulated trajectory from participants is 399Km for the head and 1053Km for both wrists. Collectively, we recorded 54M RGB images, 151.2M grayscale images, 11.7B IMUs samples and 10.8M gazes. Figure 4 shows the participant demographics w.r.t. the self-reported ethnicity, age, height and weight. The statistics is split by gender, where 48.5% participants self-identified as female, and 51.4% as male. The dataset captures 47 houses, where 31 are multi-floor. In total, there are 201 rooms and 45 gardens. In addition, we capture three locations from an campus, including 1 cafeteria with an outdoor patio, 1 multistory office building, and an open space with parking and multiple biking/hiking trails. Among the 1200 recordings, 228 are two-people sessions. Approximately 15% recordings contain outdoor activities. Among all scenarios, the highest occurrences are cooking, searching objects and improvised activity; whereas the lowest occurrences include working, locomotion, and creating a messy home. The statistics breakdown by scenarios and locations are detailed in Tab. 7 and LABEL:tab:location-breakdown in Appendix 0.C.
3.4 Privacy considerations
The Nymeria dataset was collected with rigorous privacy and ethics. We strictly follow Project Aria research guidelines for responsible innovation. Prior to data collection, consent was obtained from all participants and home owners regarding data recording and usage. Data was collected and stored in an anonymized way. The state-of-the-art algorithm EgoBlur [76] is used to blur faces and license plates in all videos, including data recorded at indoor closed environments. This de-identification process happened before further data processing and analysis.
4 Data Processing and Annotations
The data processing contains three steps. First, we use XSens software to obtain the skeleton motion, and Project Aria MPS [7] (cf. Sec. 0.B.1) for device localization, scene representation and gaze estimation. Next, motion is retarget to a parametric human model and registered into the coordinates of Aria devices via optimization. Last, data is annotated with language description.
4.1 Full-body motion capture and retargetting
We record motion at 240Hz, following the recommended procedures: 1) carefully measuring body dimensions of participants; 2) performing calibration prior to every recording; and 3) processing with the highest XSens quality with specification of single- or multi-floor of recorded activities.

### v2

@@ Abstract
We introduce Nymeria - a large-scale, diverse, richly annotated human motion dataset collected in the wild with multiple multimodal egocentric devices. The dataset comes with a) full-body ground-truth motion; b) multiple multimodal egocentric data from Project Aria devices with videos, eye tracking, IMUs and etc; and c) an third-person perspective by an additional “observer”. All devices are precisely synchronized and localized in one metric 3D world. We derive hierarchical protocol to add in-context language descriptions of human motion, from fine-grain motion narrations, to simplified atomic actions and high-level activity summarization. To the best of our knowledge, Nymeria dataset is the world’s largest human motion in the wild; first of its kind to provide synchronized and localized multi-device multimodal egocentric data; and the world’s largest motion-language dataset. It provides 300 hours of daily activities from 264 participants across 50 locations, total travelling distance over 399Km . The language descriptions contain 310.5K sentences in 8.64M words from a vocabulary size of 6545. To demonstrate the potential of the dataset we evaluate several SOTA algorithms for egocentric body tracking, motion synthesis, and action recognition.
@@
human motion egocentric multimodal dataset
@@
footnotetext: Work done during internships at Meta Reality Labs Research.
@@ 1 Introduction
The advent of AI is leading to a surge of smart glasses [30, 11, 9, 5, 6, 1, 3, 2] and other wearables. These devices not only provide seamless access to LLM-based AI assistants, but also are multimodal data-capture vehicles that provide immediate and long-term personalized context, allowing AI assistants to evolve into the next generation of human-centric contextualized AI, and unlock a new era of contextualized computing combined with AR/VR technology.
In this paradigm, the wearer’s own body motion and action provides important context. The problem is challenging to solve given insufficient self-observations from wearable devices. In reality, a critical limiting factor to advance research is data. Currently, people either use real data with limited scale, diversity and modality [69, 40, 115, 79, 108, 57, 35, 47] or simulations that lack realism and completeness [101, 46, 48, 12, 13, 60]. There are three key technical challenges to create large motion datasets.
  • Obtaining long-term ground-truth motion in the wild. There are two main motion capture (mocap) approaches. Vision-based solutions such as the ones relying on optical markers [69, 79, 57] or cameras [103, 115, 56, 63, 35] are adversely affected by line-of-sight visibility, and require a complex multi-camera-setup to cover limited range of motions inside a constrained volume. Inertial-based solutions [7, 10] suffer from dead-reckoning and thus are inferior in global positioning accuracy [91, 70].
  • Multi-device alignment. Combining multiple capture devices or ground-truth systems requires accurate temporal and spatial alignment, which can be challenging with off-the-shelf hardware that cannot be modified or lack support for universal synchronization protocols. The existing datasets work around this problem using visual cues [56, 35, 115] or audio [40]. Such approaches offer limited accuracy and reliability. In order to counter clock drift for long recordings, they can be intrusive and interrupt the natural activity. Consequently, existing datasets mostly record short motions (cf. Tab. 2).
  • Data processing and annotations. These are critical for a dataset to develop its full potential. In addition to the body motion, device localization and scene representation, we believe natural language descriptions are crucial for future research directions. Existing work provides simple descriptions or action labels without scene context [84, 36, 26], and is of significantly smaller scale compared to the text corpus for training LLMs [16, 78, 102].
Seq Qty Pts Sce Loc Pose Img IMU Gaze Traj Sent Word Voc
@@
Table 1: Highlight statistics of Nymeria dataset. We capture 1200 sequences of 300-hour daily activity from 264 people performing 20 scenarios at 50 locations with 399Km traveling distance, 260M body poses, 201.2M images, 11.7B IMUs, 10.8M gazes. The motion narrations contain 310.5K sentences in 8.64M words from 6545 vocabulary.
To fill the gap and accelerate the research, we introduce Nymeria - the world largest human motion dataset with 300 hours in-the-wild daily activities from 264 participants performing 20 scenarios from 50 indoor and outdoor locations. With avarege 15-min per recording, the data captures natural activities with spontaneous unscripted actions and authentic interactions. Nymeria is first-of-its-kind dataset recorded with multiple multimodal egocentric devices. Participants worn XSens mocap suit [7], Project Aria glasses [30] and Aria-alike wristbands to record egocentric motion, RGB, grayscale, eye tracking (ET) videos, inertial measurement units (IMUs), magnetometer, barometer and etc. Devices are synchronized with a non-intrusive hardware solution with sub-millisecond accuracy and localized into a single metric 3D leveraging Project Aria Machine Perception Service (MPS) [8]. We also developed novel algorithms to retarget XSens skeleton motion into a full parametric human model and correct the global drift with optimization. To connect human motion with natural languages, we developed a coarse-to-fine narration schema to describe in-context human motion at different granularity. With 310.5K sentences and 8.64M words from 6545 vocabulary size, Nymeria stands out as the world’s largest motion-language dataset.
dataset q/h p/M $\mu$/m pp tt/K voc pm hd 3p wd gp od gz sr mp hh
@@
Table 2: Human motion datasets by releasing date. Columns 2 to 5 show activity in hour (q/h), pose frames in millions (p/M), mean sequence duration in minute ($\mu$/m) and number of participants (pp). We then compare language narrations w.r.t. number of descriptions (tt/K) and vocabulary size (voc). The remaining columns mark following features: parametric model for motion representation (pm), egocentric head-mounted device (hd), third-person perspectives (3p), wristbands (wd), global positioning (gp), outdoor scene (od), gaze (gz), 3D scene representations (sr), multi-people scenarios (mp) and human-human interactions (hh). Note EgoExo4D [35] reports 1422h by summing per camera recording time, where the total activity is 180h, with 88.8h annotated with MSCOCO keypoints. The underline numbers are reported for the full dataset.
2 Related Works
Motion datasets – scale, multimodal, in-the-wild and perspectives.
Datasets are crucial ingredients in developing algorithms, particularly machine learning approaches. AMASS [69] is a pioneering effort in large motion dataset, which unifies multiple marker-based datasets into SMPL [65]. While AMASS provides diverse human motion, it lacks scene context. Recent works [79, 57, 47] extend the solution to include objects. Monocular [90, 81, 51, 96, 92, 33, 110, 68, 18, 63] and multi-view cameras [49, 50, 115, 56, 35] are common mocap alternatives. For monocular camera, Motion-X [63] stands out as a comprehensive large collection of whole-body motion with facial expressions and hand gestures. For multi-view setting, EgoExo4D [35] stands out as a large dataset of skilled activities. Vision-based algorithms require good line-of-sight. Consequently, they are better suited to record motion with clear body observations bounded by a volume. To capture data in the wild, mocap suit is a popular candidate [103, 70, 54, 40, 53, 59, 108, 57, 22]. To address dead-reckoning for inertial-based tracking, previous works fuse IMUs with vision [103, 70, 112], optimize motion with 3D scenes [40, 59] and limit the range of motion locally [108]. Simulation as a more scalable solution, offers valuable supplement to real data. Existing works leverage gaming engine for character animation [19], render marker-based mocap with virtual characters and scenes [12, 15, 47], simulate motion with VR [13] or by generative algorithms [60] etc. Simulations often struggle to present noise characteristics of real data, result in domain gaps. While many solutions are developed for third-person views, egocentric motion datasets remain sparse, leaving a gap to the recent advance of egocentric perception [34, 6, 30]. Existing works focus on hands with object-interaction [24, 25, 58, 93, 14], lack ground-truth [34] or parametric body motion [35], or is limited in scale, diversity, and modality [40, 12, 105, 115, 106, 56, 79, 59]. Nymeria is designed to fill the gaps with significant delta to existing datasets (cf. Tab. 2).
Motion with natural language.
Adding language descriptions to motion leads to unique perspective in motion understanding, especially given the powerful capability of large language models (LLMs). Early datasets [43, 38, 32, 41, 95, 17] offer sparse annotations as action categories or semantic labels [85]. KIT [84] is the first attempt in using complete sentences to describe locomotion. HumanML3D [36] enriches HumanAct12 [38] and AMASS [69] with multiple descriptions per motion clip. Motion-X [63] leverages narration algorithm to obtain large-scale fine-grained descriptions at sequence and frame levels. Compared to available text corpus, motion-language data is rather sparse with brief text on brief motion [22]. Ego4D [34] and EgoExo4D [35] offer large-scale descriptions of atomic actions, however, ground-truth motion is missing or without parametric representations. In our work, we provide in-context coarse-to-fine narration by annotators, with the amount of data a magnitude larger than prior works.
Egocentric body tracking, synthesis and action recognition.
Tracking one’s own body motion with wearable devices is challenging. Early approaches leverage dense body-worn IMUs [71, 42, 70]. Recent methods reduce IMUs to improve practicality [114, 48, 109, 113], assume IMU-enabled AR/VR devices [46, 45], combine IMU with cameras [120, 40, 105, 106, 112], mobile phones [73] and wristbands [59]. While sparse sensors are more practical, our work centers on constructing an offline dataset to serve “ground truth”. Consequently, we use dense IMUs to ensure high accuracy and correct global drift via optimization. Since full-body motion is ill-posed given insufficient observation from egocentric headset, motion synthesis is often used to produce plausible motion. To condition the generation, research explore sparse motion measures [108, 20, 107], headset motion [61] and eye gaze [119] and text [83, 37, 98]. The success of diffusion models lead to active develop in text-driven motion synthesis [23, 52, 99, 94, 118, 21]. Similarly LLMs inspire novel motion understanding algorithms [66, 117, 44, 31] that tightly couple motion with natural language. Our work constructs a dataset with rich hierarchical narrations to inspire further breakthroughs in the field.
3 Building Nymeria Dataset
@@
Figure 2: Capture setup. (a) A full-dressed participant. (b) The set of hardwares including Project Aria glasses, two miniAria wristbands and synchronization device. (c) The sensor suite of Project Aria and (d) the miniAria wristband.
3.1 Data collection setup
@@
Each participant wears a mocap suit, a pair of glasses, two wristbands, and a synchronization device (cf. Fig. 2. XSens MVN Link [7] is adopted for mocap, which is a tight-fit body suit wired with 17 inertial trackers and a magnetometer. MVN Link supports on-device recording, making it ideal to collect in-the-wild data. We use Project Aria glasses [30] as a lightweight headset to record multimodal data. The sensor suite includes 1 RGB camera, 2 grayscale peripheral cameras, 2 ET cameras, 2 IMUs, 1 barometer, 1 magnetometer, 7 microphones, 1 thermometer, GNSS, WiFi and BT. We repackaged the electronics and sensors of Project Aria into a new wristband device called miniAria, in order to closely resemble current AR/VR headsets and provide data to better constrain body tracking algorithms. The wristband matches Project Aria’s sensing ability, with exclusion of microphones, barometer, magnetometer and ET cameras. The supplementary provides detailed sensor configuration and recording profiles.
@@
Project Aria can record an externally provided time signal to aid synchronization. We further enable MVN Link to accept the same signal. A synchronization device is developed to supply the timestamps for all devices, which can optionally receive time from a wireless server located in radio range ($\sim$100m). This facilitates synchronizing multiple devices with sub-millisecond accuracy. The alignment between XSens and Aria is within 1 motion frame i.e. 4.2 ms. To capture multiple people simultaneously, we replicate the described setup per participant and leverage a common time server.
@@
Figure 3: Diverse scenarios by diverse people. We show different participants performing different indoor/outdoor activities at different locations. In each subfigure, we show an egocentric view on the top left, a third-person view on the bottom left, and motion rendering on the right.
3.1.3 Recording protocols.
Data collection is managed by 2-3 onsite operators. In addition to participants, a trained observer wearing Project Aria is present to record participants from third-person perspective. All people interact naturally as per activity requires, contributing to rich dynamics as opposed to staged motion. To complete each recording, participants first perform a brief mocap calibration, then gaze calibration, and finally 15-20min activity. We collect 4-8 recordings per person, where a bulk of data is captured at family houses.
@@
We define 20 scenarios (cf. examples in Fig. 1 and Fig. 3). For indoor activities, scenarios include cooking, working, entertaining, searching objects etc. For outdoor activities, scenarios include hiking, biking, dining, sports etc. To encourage natural interactions and authentic motions, participants are instructed with high-level guidelines e.g. “grab food in the cafeteria and eat on the patio”. Operators also prompt in-context actions to increase dynamics.
3.1.5 Privacy considerations.
We follow Project Aria research guideline for responsible innovation. Prior to data collection, consents were obtained from participants and home owners for recording and data usage. The SOTA de-identification algorithm EgoBlur [88] is used to blur faces and license plates for all videos.
3.2 Data processing
To process data, we first use XSens software to obtain the skeleton motion, and Project Aria MPS [8] for device localization, scene representation and gaze estimation. Then motion is retarget to a parametric human model [4] and registered into the coordinates of Aria devices via optimization.
3.2.1 Full-body mocap and retargetting.
We record motion at 240Hz, following the recommended procedures: 1) carefully measuring body dimensions of participants; 2) performing calibration prior to every recording; and 3) processing XSens with the highest quality with single- or multi-floor specification.